### v2

# DRS-OSS: Practical Diff Risk Scoring with LLMs

Abstract.
In large-scale open-source projects, hundreds of pull requests land daily, each a potential source of regressions. Diff risk scoring (DRS) can be used to estimate how likely an individual diff is to introduce a defect. This score can help prioritize reviews and tests, gate high-risk changes, and manage CI/CD capacity. Building on this idea, we present DRS-OSS, an open-source DRS tool equipped with a public API, web UI, and GitHub plugin.
DRS-OSS is a deployable, LLM-based DRS system for open-source projects built around a fine-tuned Llama 3.1 8B sequence classifier. The model consumes long-context representations that combine commit messages, structured diffs, and change metrics, and is trained on the ApacheJIT dataset. Using parameter-efficient adaptation, 4-bit QLoRA, and DeepSpeed ZeRO-3 CPU offloading, we were able to train the model with 22k-token contexts on a single 20 GB GPU, demonstrating a highly efficient training procedure.
On the ApacheJIT benchmark, DRS-OSS achieves state-of-the-art performance with F1 =0.64=0.64 and ROC-AUC =0.89=0.89. Beyond standard classification metrics, we evaluate DRS-OSS as a gating mechanism: simulations show that gating only the riskiest 30% of commits can prevent up to 86.4% of defect-inducing changes from landing. By adjusting the threshold, teams can tune the risk trade-offs during periods of high sensitivity or limited review capacity.
DRS-OSS integrates directly into developer workflows through: (i) a FastAPI gateway and LLM microservices for scalable inference, (ii) a React-based dashboard for manual diff analysis, and (iii) a GitHub App that posts risk labels and confidence scores on pull requests. The system delivers real-time, reproducible risk feedback, and is released with a full replication package including fine-tuning scripts, deployment artifacts (Salehi et al., 2025c), and source code (Salehi et al., 2025a). We additionally provide the tool’s website (Salehi et al., 2025b), along with a demonstration video showcasing the end-to-end workflow (Salehi, 2025).
Just-in-time defect prediction, Large language models, Supervised finetuning, Sequence classification, Llama
conference: ; copyright: none
1. Introduction
Open-source software (OSS) is a critical component of modern software supply chains, making its quality and reliability a matter of significant concern. However, most OSS projects lack the resources to develop or deploy advanced, industry-grade tools that help prevent defects and outages.
Diff Risk Scoring (DRS) addresses this gap by estimating, for each outgoing diff (pull request), the likelihood that a change will introduce a defect if shipped. For example, at Meta, DRS models assign a calibrated risk score to every diff and use that score to inform merge and gating decisions: low-risk diffs proceed with minimal friction, while the highest-risk changes (e.g., the top $5\%$) may be subject to short code freezes or additional reviews (Abreu et al., 2025).
Conceptually, DRS is a practical instantiation of just-in-time defect prediction (JIT-DP) in the release-deployment setting. JIT-DP reframes defect prediction at the change granularity so that risky commits can be surfaced while the author’s context is still fresh, reducing effort compared to coarser file/package models (Mockus and Weiss, 2000; Kamei et al., 2013). Classic JIT-DP formulations treat this as a binary classification problem over commits or diffs, producing a risk score and/or label for each change that can then be consumed by downstream tools, such as gating policies.
Beyond flagging risk, JIT-DP outputs can drive several software-engineering actions: (i) review prioritization: route scarce reviewer attention to the riskiest changes or assign expert reviewers first (Huang et al., 2019); (ii) test selection/prioritization: trigger a focused subset of tests for the change, reducing CI cost while preserving failure detection (Machalica et al., 2019; Srikanth et al., 2016); and (iii) policy-based gating: apply risk-aligned merge gates (e.g., flag the top-$X\%$ riskiest diffs) to reduce severe production incidents (Abreu et al., 2025).
In this work, we present DRS-OSS, a DRS tool built for open source projects. Our system is built around a fine-tuned Llama 3.1 8B sequence classifier trained on the ApacheJIT dataset (Keshavarz and Nagappan, 2022) of defect-inducing and clean commits. The model consumes long-context representations that combine commit messages, structured diffs, and change metrics as special tokens, and is trained using parameter-efficient adaptation, 4-bit QLoRA, and DeepSpeed ZeRO-3 CPU offloading, enabling 22k-token contexts while consuming only 20 GBs on GPU VRAM. The resulting risk scores can then be converted into practical gating policies for code freeze.
We make two main contributions:
  • Model: an efficiently fine-tuned Llama-based DRS classifier, evaluated on the ApacheJIT benchmark. Our results show state-of-the-art performance relative to prior JIT-DP baselines. We then use this model to analyze how DRS-style gating translates into safer code integration.
  • System: an open-source, production-oriented DRS pipeline, including an API, web dashboard, and GitHub App. This deployed system operationalizes DRS in open-source workflows, providing real-time risk labels and confidence scores on pull requests.
We note additional external interest in the tool from industry practitioners: engineering teams at Duolingo, Microsoft, and Uber have reached out regarding replicating or adapting our risk prediction pipeline for their own large-scale continuous integration environments. This feedback highlights the practical value of open, reproducible tooling for diff risk scoring in modern software development workflows.
We provide a complete replication package containing all model fine-tuning scripts (Salehi et al., 2025c) and our full implementation of the DRS-OSS service and GitHub integration (Salehi et al., 2025a).
The remainder of the paper explains DRS model training (Section 2), presents the DRS-OSS architecture and the UI (Section 5), and summarizes evaluation (Section 3) and results (Section 4).
Figure 1. End-to-end DRS-OSS architecture.
Diagram showing the GitHub App sending PR events to the Gateway/API, which calls the Llama classifier and returns a risk score rendered in the web UI.
Figure 2. Commit structuring changes the unified diff into a simpler format that can be better understood by LLMs. It also reduces the number of input tokens in the sequence.
An XML text that shows tags for COMMIT_MESSAGE, FILE, REMOVED, and ADDED enclosing various commit and diffs lines.
2. Model Training
This section describes how we train the Llama-based binary classifier that powers DRS-OSS. We choose Llama as our base model because it is pre-trained jointly on natural language and source code, making it well-suited to reason over commit messages and diffs, and because Llama-family models have already been shown to outperform classical baselines for diff risk scoring in production at Meta (Abreu et al., 2025).
We fine-tune the pretrained Llama 3.1 8B for binary sequence classification by attaching a dedicated classifier head in place of the original language-modeling head (10). Concretely, we remove the LM token-distribution layer and introduce a trainable linear projection that maps a single hidden-state vector to a scalar logit. During the forward pass, the transformer produces a hidden state for every token; we take the hidden state of the final non-pad token as a pooled representation of the entire diff, reflecting the model’s contextualized understanding of the sequence. This vector is fed through the linear projection to obtain a raw score, which is then converted to a probability via a sigmoid.
We use ApacheJIT (Keshavarz and Nagappan, 2022) as our dataset with an 80/10/10 chronological split (train/validation/test) to avoid temporal leakage, ensuring that the model is trained only on past commits and never exposed to commits from the future.
The inputs to the model are structured with special delimiters to separate commit message, code diff, and change metrics (size, entropy, files touched, etc. (Keshavarz and Nagappan, 2022)). We add these delimiters as new tokens to the model during fine-tuning so that it can learn their structural role These tokens are shown in Fig. 2. Since LLMs cannot inherently understand and compare numerical values, we feed change metrics as special tokens to the model. The special tokens are added to the tokenizer vocabulary and the embedding matrix is correspondingly resized. Placing metrics at the beginning gives the encoder early access to coarse change signals while still allowing the model to attend across message and diff content. After normalization, bucketing, and structuring, the metrics section of each sample has the following format:
[num_lines_added:] [LOW]
[num_lines_deleted:] [HIGH]
[num_files_touched:] [VERY_LOW]
[num_directories_touched:] [VERY_HIGH]
[num_subsystems_touched:] [VERY_LOW]
[change_entropy:] [MEDIUM]
[num_developers_touched_files:] [HIGH]
[time_from_last_change:] [VERY_LOW]
[num_changes_in_files:] [VERY_LOW]
[author_experience:] [LOW]
[author_recent_experience:] [MEDIUM]
[author_subsystem_experience:] [VERY_HIGH]
Class imbalance is addressed with undersampling of the majority class using a 0.7 sampling ratio during batch construction. This reduces gradient dominance from abundant negatives while keeping sufficient diversity of “clean” changes for robust thresholding later (Huang et al., 2019).
Long diffs demand aggressive efficiency measures. We adopt 4-bit weight quantization (Dettmers et al., 2023) for training to fit long-context batches on the GPU. To fine-tune efficiently, we use LoRA adapters on attention and MLP modules while keeping the frozen base weights in low precision. Sequences are truncated to a maximum length of 22,000 tokens, which covers 97.1% of commits in our corpus.
Optimization follows a memory-frugal, scalable recipe. We run DeepSpeed ZeRO-3 (Rajbhandari et al., 2021) with full parameter and optimizer offloading to CPU, enabling large effective batch sizes despite limited GPU VRAM memory In practice, this configuration consumes roughly 100 GB of host RAM during training. Gradients are accumulated across micro-batches to keep step-level statistics stable while using a per-device batch size of 1. Mixed-precision kernels are enabled where available. The model is trained for 3 epochs; on our hardware budget (a 20 GB NVIDIA A100, single device in our setup), the end-to-end run completes in roughly two weeks, including periodic validation and checkpointing. Overall, this setup allows us to fine-tune DRS-OSS with minimal computational resources while still achieving state-of-the-art results on ApacheJIT.
3. Evaluation
We evaluate and test DRS-OSS on the ApacheJIT benchmark using an 80/10/10 chronological split to avoid temporal leakage (see Section 2). We compare our model against the best JIT-DP baselines which have used ApacheJIT benchmark: JITGNN (Keshavarz and Rodríguez-Pérez, 2024), NeuroJIT (Lee et al., 2024), and CodeReviewer (Abu Talib et al., 2024).
A key part of the evaluation procedure is selecting the decision threshold, which converts the model’s predicted probability for a diff into a binary risky/safe label. Higher thresholds make the model more conservative (fewer diffs flagged as risky), while lower thresholds increase recall at the cost of more false positives. At each evaluation step during model training, we perform a sweep over candidate thresholds $\tau\in[0,1]$ and compute the F1 score on the validation split for each value. We then select the threshold that yields the best F1. When selecting the final model among multiple checkpoints, we retain both the checkpoint and its tuned threshold, and apply this fixed threshold on the held-out test split to avoid test-time tuning and ensure fairness.
4. Results
Table 1 shows the final results on the ApacheJIT benchmark. DRS-OSS achieves the strongest overall performance, including the highest F1 (0.641) and ROC-AUC (0.895). The ROC-AUC improvement over CodeReviewer (+8% relative) indicates stronger ranking ability independent of the decision threshold, while maintaining competitive precision and recall under class imbalance.
We emphasize that precision, recall, and F1 all depend on the decision threshold used to flag a diff as risky. The threshold $\tau$ maps a predicted probability $p$ to a binary label ($p\geq\tau$ implies “risky”), and varying $\tau$ inevitably trades precision against recall. These metrics further depend on the underlying class imbalance. In contrast, ROC–AUC is invariant to both the choice of threshold and the prevalence of the positive class.
Table 1. ApacheJIT test metrics at the validation-tuned best-F1 operating point. Higher is better.
Model F1 ROC-AUC Prec. Rec. Acc.
JITGNN 0.5 0.81 0.37 0.78 -
NeuroJIT 0.56 0.8 - - -
CodeReviewer 0.622 0.815 0.497 0.83 0.806
DRS-OSS 0.641 0.895 0.491 0.84 0.802
Prior works (Yang et al., 2016; Fu and Menzies, 2017) have used Recall@Top-$k\%$ for $k\!\in\!\{5,10,30\}$, i.e., the fraction of truly risky changes captured if only the riskiest $k\%$ of diffs are quarantined. Table 2 reports Recall@Top-$k\%$ for DRS-OSS. The results imply that gating only the riskiest 30% of diffs would capture 86.4% of defect-inducing commits with minimal developer disruption.
Table 2. DRS-OSS Recall@Top-$k\%$ on the test split (higher is better). Larger values at small $k$ enable stricter gates with minimal developer disruption.
Model recall@5% recall@10% recall@30%
DRS-OSS 0.315 0.503 0.864
5. System Overview
DRS-OSS delivers just-in-time risk scoring for code changes by coupling a fine-tuned Llama-3.1 8B sequence classifier with a minimal, reproducible toolchain: a Gateway/API, model services, a Web UI, and a GitHub App. The system accepts either (i) PR/commit events from GitHub or (ii) manual submissions via the Web UI, and returns a calibrated risk score and label suitable for CI gating and review triage (Fig. 1). The architecture diagram for our system is shown in Fig. 1.
Usage scenario.
  1. (1)
    Event ingress. A GitHub App subscribes to PR/commit webhooks; alternatively, a user uploads a commit message+diff in the Web UI.
  2. (2)
    Gateway/API. A FastAPI gateway routes requests to model services over REST.
  3. (3)
    Model services. Each LLM microservice validates and normalizes payloads (repo, SHA, message, diff, metadata) and applies configurations (thresholds, max diff size, timeouts) based on its settings.
  4. (4)
    Responses. The Gateway combines model outputs with policy or business logic and returns label, confidence, and optionally, explanation to the caller. For GitHub, a bot posts a PR comment; for the Web UI, it renders a result card.
Experimental explainability (off by default).
We include an experimental path that emits lightweight, risk explanations for very small diffs. This mode is not the focus of the tool demo, is disabled by default, and is omitted from evaluation due to its current limited capability; it can be toggled for exploratory use in the Web UI.
5.1. Model and API
The backend consists of a Gateway (FastAPI) and one or more Model services, each hosting a single LLM in its own Docker container. The Gateway exposes stable, documented endpoints and can be extended with database/authentication/business logic. Model services focus solely on inference and also handle input structuring of commits (Fig. 2). For relatively small commits, the API delivers inference responses in under one second.
We provide two inference modes, both implemented using Hugging Face pipeline API (Wolf et al., 2020):
  1. (1)
    Sequence-classification head. A fine-tuned Llama-3 model with a linear score head (replacing the LM head) produces calibrated probabilities for binary risk labels.
  2. (2)
    CLM-as-seq-cls. For models without an exposed classifier head (e.g., closed models), we map the first generated token to {0,1} to obtain a label; this keeps deployment flexibility while preserving a simple API.
Deployments can point the seq-cls service to (i) a full checkpoint or (ii) a PEFT adapter merged at load time. Users mount their artifacts via a host-mapped volume and configure paths through environment variables. This allows swapping models without rebuilding containers.
Each LLM service uses 4-bit weight quantization to accommodate long diffs and increase throughput. The models can be configured with mixed-precision or full-precision, sequence length limits, and quantization control.
Figure 3. Demo result on the Hive OOM-fix commit: derived label (risky/safe) and confidence for the predicted label.
A compact card showing the commit metadata and the model’s predicted risk probability p, the threshold tau, the resulting label, and whether the CI gate would allow or quarantine the change.
5.2. Web UI
Two workflows are supported via Web interface: GitHub Commit (paste a commit URL; the Gateway resolves repo/SHA and diff) and Manual Analysis (paste a commit message and a unified diff). Both return label, confidence, and optional explanation. The frontend has its own Express server and uses Vite, React, and Mantine as its tech stack.
We illustrate Web UI usage on a real change from apache/hive that fixes an out-of-memory error. Bug-fix changes, especially those touching resource management and error-handling paths, are known to be error-prone due to their complexity and cross-cutting effects (Yin et al., 2011). Fig. 3 highlights how JIT scoring is displayed in web UI.
5.3. GitHub Bot
The GitHub App integrates with repository workflows and supports both automatic and on-demand operation. The app is installed on selected repositories with read access to commit messages and code, and write access to PR comments/status checks. On pull request (opened/synchronize) or push, the app fetches the diff and commit message, calls the Gateway’s prediction endpoints, and posts the result. Reviewers can also request an explicit prediction by commenting /drs on a PR. The bot posts a compact card: commit/PR identifiers, derived label (risky/safe), confidence score, and an optional explanation if enabled.
Figure 4. GitHub bot output on a simple PR with three commits with label and confidence for each commit in the PR.
A GitHub PR comment showing a title line, a small table with label and confidence for each commit hash.
5.4. Deployment
Each LLM model runs in its own Docker container to simplify scaling and upgrades. A Docker Compose configuration brings up the Gateway, Model services, the frontend, and the GitHub App. Model services utilizes CPU offloading and eagerly load weights on startup; each service targets a single 32 GB NVIDIA V100 GPU and uses 4-bit quantization to accommodate long diffs and improve throughput.
An Nginx reverse proxy terminates TLS and routes /drs-api to the Gateway and /drs to the Web UI. A single environment file sets model IDs/paths, dtype, maximum sequence length, and quantization flags.
6. Conclusion and Community Feedback
This paper introduced DRS-OSS, a deployable JIT defect prediction system built on a fine-tuned Llama 3.1 8B model. DRS-OSS achieves state-of-the-art performance on ApacheJIT while remaining efficient enough for retraining in real-time CI workflows.
Looking forward, we plan to evaluate DRS-OSS in the context of code reviewer recommendation, where risky diffs can be routed to expert reviewers to improve review quality and reduce post-merge defects.
Finally, The tool has attracted interest from both industry practitioners and academic researchers seeking to adopt or adapt the risk prediction pipeline, underscoring the practical and research value of open, reproducible diff risk scoring tools.
Appendix A DRS-OSS Tool Walkthrough
This appendix provides a step-by-step walkthrough of the DRS-OSS tool as it would be experienced by a user during the demo. We focus on the end-to-end workflow, the Web UI, the GitHub App, and deployment/configuration details.
A.1. End-to-End Workflow
DRS-OSS delivers just-in-time risk scoring for code changes by coupling a fine-tuned Llama 3.1 8B sequence classifier with a minimal, reproducible toolchain: a Gateway/API, one or more Model Services, a Web UI, and a GitHub App.
  1. (1)
    Event ingress. A GitHub App subscribes to pull_request and push webhooks. Alternatively, a user submits a commit message and unified diff directly through the Web UI.
  2. (2)
    Gateway/API. A FastAPI gateway routes requests to the appropriate model service over REST.
  3. (3)
    Model services. Each model service runs a single LLM in its own container and exposes a prediction endpoint (e.g., /seq-cls/predict). It structures commits into DRS-OSS’s XML-like format (metrics, message, diff blocks) and returns calibrated risk scores.
  4. (4)
    Response and policy. The Gateway returns a label (e.g., risky / safe) and a confidence score.
  5. (5)
    Feedback to the user. For GitHub, a bot posts a comment or status check on the PR. For the Web UI, a result card is rendered with the risk label, confidence, and optional explanation.
A.2. DRS-OSS REST API
The DRS-OSS backend exposes a small REST API hosted at: https://worldofcode.org/drs-api. All endpoints speak HTTP/JSON and are designed to be called either by the DRS-OSS Gateway, the Web UI, or external clients (e.g., scripts, CI jobs).
A.2.1. GET /health
Simple health check to verify that the service is reachable and responsive.
GET /health curl -sS https://worldofcode.org/drs-api/health
A.2.2. POST /seq-cls/predict
Single diff risk prediction, JSON in / JSON out. The request body must contain at least a diff and a commit_message; the response returns a structured object with the predicted risk score and label.
POST /seq-cls/predict curl -sS -X POST https://worldofcode.org/drs-api/seq-cls/predict \ -H "Content-Type: application/json" \ -d ’{"diff":"diff --git ...", "commit_message":"Fix NPE"}’
A.2.3. POST /seq-cls/predict_batch
Batch prediction: an array of requests is mapped to an array of responses in the same order. This endpoint is used by the Gateway to score multiple commits from a pull request in one call.
POST /seq-cls/predict_batch curl -sS -X POST https://worldofcode.org/drs-api/seq-cls/predict_batch \ -H "Content-Type: application/json" \ -d ’[{"diff":"...","commit_message":"..."}, {"diff":"...","commit_message":"..."}]’
A.2.4. POST /seq-cls/predict_by_sha
Prediction by repository and commit SHA. The backend looks up the diff for the given commit and returns the same JSON schema as /seq-cls/predict.
POST /seq-cls/predict_by_sha curl -sS -X POST https://worldofcode.org/drs-api/seq-cls/predict_by_sha \ -H "Content-Type: application/json" \ -d ’{"owner_repo":"facebook/react", "commit_sha":"abc123..."}’
A.2.5. POST /clm/predict
Causal language model (CLM) endpoint: JSON in, raw text out. This is used for experimental natural-language explanations of a risk score.
POST /clm/predict curl -sS -X POST https://worldofcode.org/drs-api/clm/predict \ -H "Content-Type: application/json" \ -d ’{"diff":"...","commit_message":"..."}’
A.2.6. POST /clm/predict_by_sha
CLM prediction by repository and commit SHA. Returns a raw text response (i.e., a short rationale) for the given commit.
POST /clm/predict_by_sha curl -sS -X POST https://worldofcode.org/drs-api/clm/predict_by_sha \ -H "Content-Type: application/json" \ -d ’{"owner_repo":"facebook/react", "commit_sha":"abc123..."}’
A.3. Web UI Walkthrough
The Web UI is intended for manual analysis, debugging, and demonstration. It is built with React, Vite, and Mantine, and communicates with the Gateway over REST.
A.3.1. Landing Page and Input Modes
On the landing page (Figure 5), users can choose between two main workflows:
  • GitHub Commit / PR mode. The user pastes a GitHub URL (commit or pull request). The microservices resolve the repository and SHA(s), fetch the diff, and forward a normalized payload to the model service.
  • Manual Analysis mode. The user directly pastes a commit message and unified diff into a text area. This mode is independent of any Git hosting service and is useful for local experiments or offline examples.
Figure 5. DRS-OSS Web UI landing page with GitHub and manual analysis modes.
A.3.2. Submitting a Commit and Viewing Results
Once a commit or PR is submitted:
  1. (1)
    The UI shows a pending status while the Gateway calls the model service.
  2. (2)
    When the response arrives, a result card displays:
    • the derived label (risky or safe);
    • the model’s confidence (probability);
    • optionally, an explanation snippet for small diffs (if enabled).
A typical prediction view is shown in Figure 6.
Figure 6. Web UI result card for a real-world commit, showing label and confidence.
A.3.3. Experimental Explanations
For very small diffs, an experimental explanation path can be enabled:
  • The Gateway issues a secondary request to the explanation endpoint.
  • The UI displays a short natural-language rationale (e.g., “touches error-handling and memory management code”) beneath the main result card.
This feature is off by default and at this point it is only experimental. An example explanation view is shown in Figure 7.
Figure 7. Optional explanation view for a small diff, showing a short natural-language rationale.
A.4. GitHub App Walkthrough
The GitHub App integrates DRS-OSS directly into repository workflows and supports both automatic and on-demand operation.
A.4.1. Automatic Predictions on PR Events
When installed on a repository, the app listens to pull_request events (opened, synchronize):
  1. (1)
    On a new or updated PR, the app fetches the associated commits, messages, and diffs.
  2. (2)
    It sends a prediction request to the DRS-OSS Gateway for each commit.
  3. (3)
    After inference, the app posts a comment listing per-commit labels and confidences (Figure 8).
Figure 8. GitHub App comment summarizing per-commit labels and confidence scores on a pull request.
A.4.2. On-Demand Predictions via Commands
Reviewers can also request predictions explicitly by posting a command (e.g., /drs) as a comment on a PR:
  • The bot parses the command, fetches the current diffs, and calls the Gateway.
  • It then replies with an updated result card, allowing reviewers to re-check risk after code changes or rebases.
A.5. Configuration and Deployment
DRS-OSS is designed to be reproducible and easy to deploy using Docker and Docker Compose.
A.5.1. Services and Configuration
A typical demo deployment includes the following services:
  • Gateway (FastAPI). Exposes REST endpoints (e.g., /api/v1/drs/predict) and handles routing, and policy.
  • Model services. One or more containers each hosting a single LLM checkpoint, using 4-bit quantization to accommodate long diffs on a single GPU.
  • Web UI. A React/Vite frontend served via an Express server.
  • GitHub App (optional). A small service that handles webhook ingress and outgoing API calls to GitHub.
Configuration is controlled via a single .env file, which specifies:
  • model identifiers or local paths (e.g., SEQCLS_DRSLLM_MODEL_ID, SEQCLS_BASE_MODEL_PATH);
  • sequence length and quantization flags (e.g., DRSLLM_MAX_LENGTH, DRSLLM_LOAD_IN_4BIT);
  • gateway URLs, authentication tokens, and GitHub App credentials.